James can stuff 1000 envelopes in 5 hours. Susan can stuff 1000 envelopes in 3 hours. How long will it take to stuff 1000 envelopes if they worked together?

answer:it will take 1 7/8 hours to stuff 1,000 envelopes working together
need a break down of how my book got 1 7/8 hours to stuff 1,000 envelopes together

Found 2 solutions by ptaylor, themathprof:
Answer by ptaylor(2052)

Let x=number of hours it takes with both working together
James stuffs at the rate of 1000/5=200 envelopes per hour
Susan stuffs at the rate of 1000/3=333.333333 =333 1/3 envelopes per hour
Together they stuff at the rate of 200+333.3333=533.3333=533 1/3 envelopes per hour
So, our equation to solve is:
(533 1/3)*x=1000 divide each side by 533 1/3
x=1.875 hours or 1 7/8 hours--------ans
Now we can stay with fractions:
x=1000/(533 1/3)=1000/(1600/3)=1000*(3/1600)=3000/1600=30/16=15/8=1 7/8 hr
Hope this helps----ptaylor

Answer by themathprof(21)

James can stuff 1000 envelopes in 5 hours. Susan can stuff 1000 envelopes in 3 hours. How long will it take to stuff 1000 envelopes if they worked together?
The solution does not depend on the number 1000. It would be solved the same way even if 10 letters were stuffed, or 100,000 letters were stuffed.
Think of it as completing a job.
James can complete a job 5 hours. Susan can complete the same job in 3 hours. How long will it take to complete the job if they worked together?
Let x=amt of hours each will work. Since they're working together, each of their times =x
James needs 5 hrs by himself. So each hour he completes 1/5.
Susan needs 3 hours by herself, so each hour she completes 1/3
But each don't work 1 hour they work x hours. So,...
James completes x/5
Susan completes x/3
Together they complete x/5+x/3
The whole job is 1
x/5+x/3=1
Multiply each term by the LCM=15 to clear fractions
15x/5 +15x/3=15
3x+5x=15
8x=15
x=15/8
x= 1 7/8 hours
See more word problems like this
worked out step by step in my FREE
math911 tutorials. Download, install,
activate, ALL FREE from
www.math911.com
Tell your friends, (and teacher!)
Professor Martin Weissman